HIGH RESOLUTION FOURIER TRANSFORM EMISSION SPECTROSCOPY OF THE Ã^{2}Σ^{+} −X̃^{2}Π TRAN-SITION OF THE BrCN^{+} ION

\underline{YOSHIHIRO NAKASHIMA}, TOMOKI OGAWA, MAKI MATSUO, and KEIICHI TANAKA, Department of Chemistry, Faculty of Science, Kyushu University 33, Hakozaki, Higashiku, Fukuoka 812-8581, Japan.

The BrCN^{+} ion was produced by Penning ionization of BrCN with metastable He^{∗}, and the near-infrared emission in the 690 - 870 nm region was measured by Fourier transform spectrometer. Thirteen vibronic bands of the Ã^{2}Σ^{+} − X̃^{2}Π transition of ^{79}BrCN^{+} and ^{81}BrCN^{+} were observed. The rotational analysis was performed for the Ã^{2}Σ^{+}(000) − X̃^{2}Π_{Ω}(000) and Ã^{2}Σ^{+}(000) − X̃^{2}Π_{Ω}(010) transitions, both for the Ω = 3/2 and 1/2 spin components.

Molecular constants, including the effective rotational constant, centrifugal distortion constant, Λ-type doubling constant in the X̃^{2}Π_{1/2} state, and spin-rotation interaction constant in the Ã^{2}Σ^{+} state, were determined from the observed spectrum. Spin-orbit interaction constants for the X̃^{2}Π ground state were determined to be −1476.4669(48) and −1476.4841(60) cm^{−1}, respectively, for ^{79}BrCN^{+} and ^{81}BrCN^{+}. The r_{0}-structures for the X̃^{2}Π and Ã^{2}Σ^{+} states of BrCN^{+} were derived to be compared with that for the X̃^{1}Σ^{+} state of BrCN. The geometrical change of the BrCN^{+} ion from the BrCN molecule was turned out to be small.

Due to the Renner-Teller effect, the X̃^{2}Π(010) state was split into four components, μ^{2}Σ, κ^{2}Σ, and ^{2}Δ_{P} (P = 5/2 and 3/2), and the rotational analysis was performed both for the Ã^{2}Σ^{+} − μ^{2}Σ and Ã^{2}Σ^{+} − κ^{2}Σ transitions. For the μ^{2}Σ and κ^{2}Σ vibronic states, Ω-type doubling constants were determined as well as the rotational constants and centrifugal distortion constants. Renner parameter ϵ for the X̃^{2}Π state was determined to be −0.185 from the Ω-type doubling constants for both the isotopic species.